White light is normally incident on a diffraction
Answer #1

First we need to find the Angle ,\theta=sin-1(m\lambda/d)

Where m=1

Wavelength,\lambda=700 nm for red and 400 nm for violet

d=6500 lines /cm

So \thetaR=sin-1(700/6500)=6.18

\thetaV=sin-1(400/6500)=3.52

So the difference in angle is ,d\theta=2.65